Calculate pI for the titration of 50.0 mL of 0.0250 M KI after adding 25.0 ml of
0.0500 M AgNO3.
1
Expert's answer
2020-06-23T02:16:35-0400

Moles of KI =(50/1000) × 0.025 mole.

Moles of AgNO3 = ( 25/1000) × 0.05 mole .

KI + AgNO3 ------ AgI + KNO3 .

After Reaction of KI and AgNO3 Product formed are AgI and KNO3 moles of AgI and KNO3 formed are 1.25 × 10 -3 moles. And both KI and AgNO3 are totally consumed .

Concentration of AgI = (1.25 / 75 ) mole / litre .

So , Concentration of I = (1.25 / 75 ) M.


So , PI = -log l = -log 1.25/75 = 1.78 . Ans...